OpenStudy (anonymous):

Given that f(x) = 2x + 1 and g(x) = the quantity of 3x minus 1, divided by 2, solve for g(f(3)). Please help I'm so bad at algebra 2

OpenStudy (anonymous):

g(x) = (3x - 1) / 2 f(x) = 2x + 1 start inside out for g(f(3)) f(3) = 2(3) + 1 f(3) = 7 g(f(3)) = (3(7) - 1) / 2 g(f(3)) = (21 - 1) / 2 g(f(3)) = 10

OpenStudy (anonymous):

Thank you so much!!!!